# S3E: A Mulit-Robot Multimodal Dataset for Collaborative SLAM

### v6

@@ III-D Trajectory Paradigms
Concentric Circles: In this paradigm, multiple unmanned platforms move in a coordinated formation, maintaining interconnectivity and observing consistent environmental data. This setup enables continuous loop detection among plat- forms, enhancing the system’s ability to perform comprehensive area coverage and exploration tasks such as environmental monitoring and map construction. The real-time sharing of perceptual information and pose estimation results among platforms bolsters the system’s positioning and mapping accuracy. However, the high similarity of observed data may complicate the detection and management of dynamic environmental changes.
Intersecting Circles: This paradigm is characterized by individual unmanned platforms operating in distinct zones, exchanging perceptual information upon intersection. Communication is established only at these points, with partial environmental information overlap. The exchange of historical data for loop closure detection at intersections, along with independent loop closure capabilities of each platform, expands the exploration range and improves mission efficiency. The division of search areas reduces communication and coordination demands within the system. However, limited interaction between platforms may hinder the achievement of globally consistent mapping and positioning.
@@
Intersection Curve: Unmanned platforms initiate from diverse starting points, intersect at designated locations, and converge at a common endpoint. Communication and loop closure detection occur only upon meeting, with partial environmental information overlap. This paradigm is adept for large-area distributed exploration and mapping, facilitating regular information exchange and loop closure detection to minimize cumulative errors. The independent exploration by each platform enhances task efficiency and coverage. However, higher navigation and control capabilities are required for planning suitable intersection points and paths.
Rays: In this paradigm, platforms depart from various starting points, moving independently until converging at a shared endpoint. Communication is restricted to the final convergence, with minimal environmental information overlap. The platforms can exchange historical information for loop closure detection at convergence, but lack individual loop closure detection capabilities. This paradigm is suitable for large-scale independent exploration and mapping tasks, such as post-disaster searches and mine explorations. It emphasizes the autonomous execution of tasks by each platform, relying on its perceptual and computational resources. The final convergence allows for historical information exchange and local map optimization. However, the scarcity of information exchange poses challenges for globally consistent mapping and positioning, necessitating advanced autonomous navigation and decision-making.
@@
Table IV: Analysis of S3E dataset.
@@ III-E Dataset Analysis
To align with the collaborative mission scenarios and to reflect the nuances of the four trajectory paradigms, we have strategically chosen a variety of complex and representative scenes across different areas within Sun Yat-sen University’s Guangzhou East Campus for our data collection endeavors. The mobile platforms, comprising three tele-operated robots, are under human control to ensure safety during operations.
For an exhaustive overview of the trajectory paradigms applied, the lengths of the trajectories, and the types of sensors utilized for each sequence, please refer to LABEL:Tab:Analysis-of-S3E. The dataset encompasses a diverse range of environments, including squares, libraries, and six additional regions, comprising a total of 13 outdoor and 5 indoor sequences. The aggregated size of all 18 sequences is over 263.2 gigabytes, with a combined duration of more than 475.2 minutes and an extensive operating distance surpassing 28.1 kilometers. On average, the sequences last for 527.9 seconds each, which validates the dataset’s adequacy for extensive, long-term assessments of C-SLAM technologies.
Furthermore, the dataset encompasses a variety of trajectories that adhere to the paradigms outlined in Section LABEL:Subsec:Trajectories. This diversity is instrumental for rigorous testing and comparative analysis of different C-SLAM methodologies under various interactive conditions, such as intra-robot and inter-robot loop closures.
@@
Dormitory: Characterized by high pedestrian traffic and the presence of dynamic objects such as pedestrians and bicycles, dormitory areas challenge the perception and tracking abilities of unmanned platforms. Their regular architectural layouts offer an ideal setting for evaluating the precision and consistency of C-SLAM algorithms across multiple platforms.
Campus Road: Serving as critical connectors, campus roads are marked by long distances and expansive views. They test the endurance and large-scale exploration capabilities of unmanned platforms. Extensive data collection along these roads, including long-distance and multi-cycle datasets, provides a robust foundation for assessing the stability, accuracy, and efficiency of C-SLAM algorithms over extended operations.
Playground: As open areas with fewer obstructions, sports fields challenge feature extraction, registration, and optimization processes. Data collected at various times of day, including day and night, evaluates the adaptability of C-SLAM algorithms to different lighting conditions. Additionally, data from rapid motion scenarios assesses the algorithms’ performance under significant motion.
Laboratory: Indoor environments like laboratories, with their confined spaces, complex layouts, and rich semantic content, challenge navigation, obstacle avoidance, and semantic mapping. The presence of various instruments and furniture makes laboratories suitable for testing the advanced scene understanding and mapping capabilities of C-SLAM algorithms.
Teaching Building and Tunnel: These areas, with their severe perceptual aliasing, pose significant challenges due to poor lighting in tunnels and similar geometric structures in corridors, which can lead to errors in data association. They test the robustness of C-SLAM algorithms in maintaining accurate positioning and mapping.
Furthermore, our dataset includes other distinctive regions such as libraries and squares, which contribute additional environmental features and perceptual challenges to the dataset’s diversity and comprehensiveness. LABEL:Fig:outdoor_tracks displays the various trajectory paradigms captured in outdoor environments within the S3E dataset. Our dataset encompasses a diverse array of challenging environments that are likely to be faced by C-SLAM algorithms in real-world deployments. These include dynamic settings with moving objects, prolonged operational durations, scenarios with perceptual aliasing, indoor environments, and situations characterized by substantial motion. The inclusion of such a wide range of conditions lays a solid groundwork for an in-depth evaluation of the performance metrics, adaptability to different conditions, and overall robustness of C-SLAM algorithms. This thorough vetting process is essential for advancing the state-of-the-art in collaborative robotic navigation and mapping technologies.
Table V: Descriptions of ROS2 topics included in our dataset. Note that agent_id stands for Alpha, Blob, and Carol.
@@ III-F Dataset Format
We employ the ROS2 [31] bag format for storing the sensor data, which is a widely-recognized standard in the robotics community. This format allows for efficient data management and playback, enabling users to simulate and analyze the dataset in various software environments. For the purpose of simplifying data replay and ensuring seamless access to the synchronized sensor streams from our robotic platforms, we have implemented a consolidation strategy. Specifically, we merge the sensor data collected from the Alpha, Blob, and Carol robots during the same operational sequence into a single ROS2 bag file.
Data Organization: As depicted in LABEL:Fig:data_organization , our dataset features a well-organized directory structure. The calibration parameters are documented in YAML files and stored in the dedicated Calibration directory. Each sequence is accompanied by essential data files, including .db3 file that house the primary sensor measurements, and metadata.yaml file that provide essential information about the sequence, such as sensor settings and capture conditions. Furthermore, to facilitate comprehensive performance evaluation, we have included auxiliary files such as <agent_id>_gt.txt, which contains the ground truth data.
@@ IV Experiments
Table VI: ATE $[m]$ for Single SLAM and C-SLAM in the S3Ev1.0 outdoor environment without UWB measurement. $\alpha$, $\beta$, and $\gamma$ stand for Alpha, Bob, and Carol respectively. ✗ fails to initialize or track frames. If inter-loop closures detection fails, we mark it “Failed”.
@@
Table VII: ATE $[m]$ for Single SLAM and C-SLAM in the S3Ev1.0 indoor environment without UWB measurement. $\alpha$, $\beta$, and $\gamma$ stand for Alpha, Bob, and Carol respectively.
@@
Table VIII: ATE $[m]$ for C-SLAM in the S3Ev2.0 outdoor environment with UWB measurement. $\alpha$, $\beta$, and $\gamma$ stand for Alpha, Bob, and Carol respectively.
@@ IV-A Baselines
[…]
[autoplay,loop,width=]10videos/360/teaching_building/000096
(a) Teaching_building_1.
[…]
[autoplay,loop,width=]10videos/360/square_1/000197
(b) Square_1.
Figure 9: The animations of S3E dataset. For an optimal viewing experience, we recommend using Adobe Reader / KDE Okular , as these applications provide the necessary support for smooth playback and clear visualization of the animated sequences.
@@ IV-B Results
The results from our experiments on the S3E dataset underscore the importance of continuous innovation in C-SLAM algorithms. The varying performance across different scenarios and the impact of collaborative data sharing highlight the need for further research. The qualitative results presented in LABEL:Fig:qualitative_result_2 and LABEL:Fig:qualitative_results_3, along with the animations in LABEL:Fig:animations, visually complement the performance metrics, providing a deeper understanding of the C-SLAM systems’ behavior in diverse environments.
@@ V Conclusion
In conclusion, the S3E dataset represents a significant advancement in the field of Collaborative Simultaneous Localization and Mapping (C-SLAM). Its expansive and multimodal nature, coupled with the introduction of four distinct trajectory paradigms, provides an unparalleled platform for evaluating C-SLAM systems across a multitude of real-world scenarios. Captured across various indoor and outdoor environments, the dataset includes meticulously synchronized streams from a range of sensors, offering an unprecedented level of detail and complexity. Our experiments utilizing this dataset have highlighted the improved robustness of C-SLAM systems, especially in their ability to handle inter-loop closures. The incorporation of UWB measurements has further enhanced the accuracy and reliability of localization within these systems. The S3E dataset serves not only as a benchmark for current C-SLAM methodologies but also as a catalyst for future innovations in multi-agent robotic navigation and mapping. We anticipate that this dataset will inspire further research and development, pushing the boundaries of what is possible in collaborative robotic systems.

### v7

@@ III-D Trajectory Paradigms
Concentric Circles: This paradigm involves coordinated unmanned platforms that stay connected and collect similar environmental data. This allows for efficient area coverage and tasks like environmental monitoring and mapping. Real-time data sharing improves accuracy, but the similarity of observations can make it challenging to detect and adapt to rapid environmental changes.
Intersecting Circles: In this paradigm, unmanned platforms work separately in different areas and only share data when they intersect. This approach reduces the need for constant communication and coordination but may limit the consistency of overall mapping and positioning due to less frequent data exchange.
@@
Intersection Curve: This paradigm involves unmanned platforms starting from different locations, meeting at specific points, and eventually converging at a final destination. They communicate and share data only during these meetings, which helps in large-scale exploration and mapping by reducing cumulative errors. Each platform’s independent operation boosts efficiency and coverage. However, this approach requires advanced navigation for planning effective meeting points and routes.
Rays: In this paradigm, autonomous platforms begin at various starting points, work independently, and come together at a common endpoint for communication and data exchange. This method is effective for large-scale, independent operations like disaster recovery and mining exploration, where each platform operates on its own. The final convergence helps in sharing historical data and optimizing local maps, but the limited interaction can affect the overall consistency of mapping and positioning, thus requiring advanced autonomous navigation and decision-making skills.
@@
Table III: Analysis of S3E dataset.
@@ III-E Dataset Analysis
To test collaborative mission scenarios and trajectory paradigms, we conducted data collection at Sun Yat-sen University’s Guangzhou East Campus using three tele-operated robots under human control for safety. The dataset, detailed in LABEL:Tab:Analysis-of-S3E, includes diverse environments like squares and libraries, with 13 outdoor and 5 indoor sequences. It totals over 263.2 GB, 475.2 minutes of footage, and covers more than 28.1 km. With an average sequence length of 527.9 seconds, it’s suitable for long-term C-SLAM technology evaluations.
@@
Dormitory: High pedestrian and bicycle traffic test perception and tracking. Regular layouts are ideal for assessing C-SLAM precision and consistency across platforms.
Campus Road: Long distances and expansive views evaluate endurance and large-scale exploration. Long-distance and multi-cycle data assess stability, accuracy, and efficiency over extended operations.
Playground: Open spaces with fewer obstructions challenge feature extraction and optimization. Data collected at different times of day test adaptability to lighting conditions and performance under rapid motion.
Laboratory: Confined spaces with complex layouts and rich semantic content test advanced scene understanding and mapping capabilities.
Teaching Building and Tunnel: Poor lighting and similar geometric structures challenge robustness in maintaining accurate positioning and mapping, especially in tunnels and corridors.
Our dataset features unique areas like libraries and squares, adding environmental variety and perceptual challenges. LABEL:Fig:outdoor_tracks shows the trajectory paradigms in outdoor settings within the S3E dataset. The dataset covers a range of challenging environments that C-SLAM algorithms may encounter, including dynamic objects, long operation times, perceptual aliasing, indoor settings, and significant motion. This diversity is crucial for evaluating C-SLAM performance, adaptability, and robustness, which are key for advancing collaborative robotic navigation and mapping technologies.
Table IV: Descriptions of ROS2 topics included in our dataset. Note that agent_id stands for Alpha, Blob, and Carol.
@@ III-F Dataset Format
Our research utilizes the ROS2 [14] bag format for sensor data storage, a standard in robotics known for efficient data management and playback. To simplify data replay and ensure easy access to synchronized sensor streams from our Alpha, Blob, and Carol robots, we’ve implemented a strategy to merge data from the same operational sequence into one ROS2 bag file.
Data Organization: Our dataset, as shown in LABEL:Fig:data_organization , is meticulously organized with a clear directory structure. Calibration parameters are detailed in YAML files within a specific Calibration folder. Each data sequence is complete with a .db3 file for primary sensor measurements and a metadata.yaml file outlining sequence specifics like sensor configurations and capture conditions. Additionally, to aid in thorough performance assessments, we’ve included auxiliary files like <agent_id>_gt.txt, which contain ground truth data.
@@ IV Experiments
Table V: ATE $[m]$ for Single SLAM and C-SLAM in the S3Ev1.0 outdoor environment without UWB measurement. $\alpha$, $\beta$, and $\gamma$ stand for Alpha, Bob, and Carol respectively. ✗ fails to initialize or track frames. If inter-loop closures detection fails, we mark it “Failed”.
@@
Table VI: ATE $[m]$ for Single SLAM and C-SLAM in the S3Ev1.0 indoor environment without UWB measurement. $\alpha$, $\beta$, and $\gamma$ stand for Alpha, Bob, and Carol respectively.
@@
Table VII: ATE $[m]$ for C-SLAM in the S3Ev2.0 outdoor environment with UWB measurement. $\alpha$, $\beta$, and $\gamma$ stand for Alpha, Bob, and Carol respectively.
@@ IV-B Results
The results from our experiments on the S3E dataset underscore the importance of continuous innovation in C-SLAM algorithms. The varying performance across different scenarios and the impact of collaborative data sharing highlight the need for further research. The qualitative results presented in LABEL:Fig:qualitative_result_2 and LABEL:Fig:qualitative_results_3, visually complement the performance metrics, providing a deeper understanding of the C-SLAM systems’ behavior in diverse environments.
@@ V Conclusion
The S3E dataset provides a comprehensive and multifaceted platform for evaluating C-SLAM systems under a variety of real-world conditions. It features four distinct trajectory paradigms for broad scenario evaluation. Captured across diverse indoor and outdoor environments, the dataset includes meticulously synchronized sensor data, offering high detail and complexity. Our experiments using this dataset have highlighted the improved robustness of C-SLAM systems, especially in handling inter-loop closures. The addition of UWB measurements has significantly enhanced localization accuracy and reliability. The S3E dataset establishes a new benchmark for C-SLAM, fostering further research and innovation in multi-agent robotic navigation and mapping, and advancing the field.
Acknowledgments
This research greatly benefited from the guidance and expertise of many contributors. We extend our profound gratitude to our colleagues, Yizhen Yin and Haoxin Zhang from Sun Yat-sen University, for significantly improving our dataset’s quality and applicability, especially in data collection. Special acknowledgment goes to Prof. Tao Jiang and his student Yudu Jiao from Chongqing University for evaluating the single-agent SLAM algorithms on our dataset.